Birds use color to select and avoid certain types of food. The authors of a particular article studied the pecking behavior of 1-day-old bobwhites. In an area painted white, they inserted four pins with different colored heads. The color of the pin chosen on the bird's first peck was noted for each of 35 bobwhites, resulting in the accompanying table.

Color First
Peck
Frequency
Blue 15
Green 9
Yellow 7
Red 4

Do these data provide evidence of a color preference (i.e. are the colors not all the same)? Test using α = 0.01. (Assume that the sample is representative of all 1-day-old bobwhites.)

Let p1, p2, p3, and p4 be the proportions of bobwhites whose first peck is each of the four colors.

State the appropriate null and alternative hypotheses.

H0: p1 = p2 = p3 = p4 = 0.5
Ha: H0 is not true. H0: p1 = p2 = p3 = p4 = 35
Ha: H0 is not true.     H0: p1 = p2 = p3 = p4 = 0.01
Ha: H0 is not true. H0: p1 = p2 = p3 = p4 = 9
Ha: H0 is not true. H0: p1 = p2 = p3 = p4 = 0.25
Ha: H0 is not true.

Find the test statistic and critical value. Round your test statistic to three decimal places and your critical value to two decimal places.)

X2 = critical value =

State the conclusion in the problem context.

Fail to reject H0. There is convincing evidence that there is a color preference. Fail to reject H0. There is not convincing evidence that there is a color preference.     Reject H0. There is convincing evidence that there is a color preference. Reject H0. There is not convincing evidence that there is a color preference.
Answer #1

The null and alternative hypotheses are H: P = P2 = p; = P4 = 0.25 H:H, is not true E = np = 35 = 35(0.25) = 8.75 1-1 x = 2 (The value of the test statistic is x2 = 7.400. The critical value is 11.35. Fail to reject H. There is not convincing evidenc